# GPU-accelerated LISA parameter estimation with full time domain response

### v2

@@ Abstract
We conduct the first full Bayesian inference analysis for LISA parameter estimation incorporating the effects of subdominant harmonics and spin-precession through a full time domain response. The substantial computational demands of using time domain waveforms for LISA are significantly mitigated by implementing a novel Python version of the IMRPhenomT family of waveform models and the LISA response with GPU acceleration. This time domain response alleviates the theoretical necessity of developing specific transfer functions to approximate the LISA response in the Fourier domain for each specific type of system and allows for the use of unequal arms configurations and realistic LISA orbits. Our analysis includes a series of zero-noise injections for a Massive Black Hole Binary with aligned and precessing spins. We investigate the impact of including subdominant harmonics, compare equal and unequal arm configurations, and analyze different Time-Delay-Interferometry (TDI) configurations. We utilize full and uniform priors, with a lower frequency cutoff of 0.1 mHz, and a signal duration of approximately two months, sampled every 5 seconds. The sampler is initialized based on Fisher’s estimates. Our results demonstrate LISA’s capability to measure the two spin magnitudes and the primary spin tilt angle, alongside sky localization, with percent-level precision, while component masses are determined with sub-percent accuracy.
@@ I Introduction
The Laser Interferometer Space Antenna (LISA) mission Colpi et al. (2024) was officially adopted by ESA and NASA in January 2024, with a scheduled launch by 2035. LISA will deploy three identical spacecraft into space to form the first space-based gravitational wave (GW) interferometer. Operating in space allows LISA to bypass many of the noise limitations that terrestrial detectors face Abbott et al. (2020), enabling the observation of a new class of GW sources. LISA will observe in the millihertz frequency range, detecting signals from a broad spectrum of sources, from stellar-mass compact objects to supermassive black holes, providing us with unprecedented insights into fundamental physics. LISA promises to be a transformative mission, set to revolutionize our understanding of the universe and GW astronomy.
@@
Moreover, LISA’s ability to observe a wide range of masses requires waveform models that encompass an expanded parameter space. At the same time, the high signal-to-noise ratios expected from LISA events will demand highly accurate models capable of capturing subtle physical effects that are less noticeable with current detectors. Key among these are precession and eccentricity, which are essential to model as they provide crucial information on the astrophysical formation and evolutionary pathways of compact binaries Breivik et al. (2016); Steinle and Kesden (2021); Zevin et al. (2021); Fumagalli and Gerosa (2023); Fumagalli et al. (2024). Developing efficient parameter estimation techniques that incorporate both precession and eccentricity will be of paramount importance for fully exploiting LISA’s scientific potential.
@@
Parameter estimation for gravitational waves is typically performed in the Fourier domain due to the easier characterization of the noise and the efficiency and sensitivity of the matched filtering technique Allen et al. (2003, 2012), where the complicated convolutions and correlations between templates and data become a simple multiplication in the Fourier domain. Consequently, waveform models formulated directly in the Fourier domain Pratten et al. (2020); García-Quirós et al. (2020); Pratten et al. (2021); Cotesta et al. (2020) have gained significant popularity in GW data analysis Abbott et al. (2023); Colleoni et al. (2021); Mateu-Lucena et al. (2022). By avoiding the need for a discrete Fourier transform of each template, these models save valuable computational resources. In contrast, time domain waveforms require fine sampling on a uniformly spaced time grid to facilitate the application of the Fast Fourier Transform (FFT) algorithm. This constraint makes time domain waveforms particularly inefficient when dealing with long-duration signals, as the sampling grid must encompass the entire signal duration, even if one is interested only in a reduced frequency range. This also prevents the use of custom non-uniform frequency grids, which could be used to optimize performance Vinciguerra et al. (2017); García-Quirós et al. (2020); Canizares et al. (2015).
@@
However, from the modelling perspective, the time domain is often preferred because it naturally represents the temporal evolution of physical systems. Numerical Relativity (NR), Effective One Body formalism (EOB) Buonanno and Damour (1999); Damour and Nagar (2009) or post-Newtonian (PN) theory Will (2011); Blanchet (2014) solve the dynamic equations of motion in the time domain, yielding time domain waveforms that accurately describe the intricate behaviour of binary systems. In this domain, it is easier to incorporate physical processes and time-varying interactions, making it easier to asses how physical details like spin precession or eccentricity affect the waveform’s shape. Moreover, certain phenomena like displacement memory, a permanent change in spacetime following the passage of gravitational waves, are more naturally described and understood in the time domain Rosselló-Sastre et al. (2024). As LISA aims to explore new physics through these effects, time domain models will be crucial for maximizing the scientific return from the mission, enabling precise waveform modelling that is harder to replicate when relying solely on frequency domain representations.
@@
Apart from the challenges in waveform modelling, LISA introduces the additional complexity of computing a dynamic detector response. For current ground-based detectors, signals are typically short-lived, allowing the detector to be approximated as static. In these cases, the detector response is simply a linear combination of the two gravitational wave polarizations, with the response factors depending only on the sky position and polarization angle, but not on time. However, LISA’s sources can remain in-band for extended durations, from days to months or even years Colpi et al. (2024), causing the signal to be modulated by the orbital motion of the LISA constellation. Moreover, the laser frequency noise is several orders of magnitude larger than the GW signal LISA aims to detect. This noise is exacerbated when the LISA arms have unequal length. In a realistic scenario, the LISA spacecraft will not maintain a perfect equilateral triangular formation, leading to time-varying arm lengths and complicating gravitational wave detection.
Fortunately, the Time-Delay Interferometry (TDI) technique Tinto and Dhurandhar (2014) mitigates this issue by combining signals from different arms in a carefully timed manner, effectively cancelling out laser frequency noise by several orders of magnitude. TDI channels reconstruct gravitational wave signals while suppressing noise, thereby enhancing the sensitivity of the detector to gravitational waves.
@@
The LISA arm response and TDI are better understood and described in the time domain. In this domain, they can be applied to the GW $h_{+,\times}(t)$ polarizations exhibiting any physical effect: higher harmonics, precession, eccentricity, or other complex features. However, applying these operations in the Fourier domain is more challenging and requires the development of theoretical transfer functions Marsat and Baker (2018); Marsat et al. (2021) to approximate the LISA response. These transfer functions rely on the assumption that a unique time-frequency correspondence exists in the gravitational wave signal. This requirement for a monotonic frequency and amplitude evolution complicates this Fourier domain approach, particularly for complex signals. Marsat et al. Marsat et al. (2021) addressed this for subdominant harmonics signals in aligned spin systems by developing a specific transfer function for each harmonic. Later, Pratten et al. Pratten et al. (2023) extended this to precessing systems by introducing the multiple-scale analysis to provide first-order corrections to the transfer function for each harmonic in precessing systems. These works marked significant progress, as they were the first to perform full Bayesian inference using complete inspiral-merger-ringdown (IMR) Fourier-domain waveforms for LISA. In contrast, Garg et al. Garg et al. (2024) applied this formalism to eccentric systems, using a Fourier-domain inspiral-only approximant without higher harmonics. So far there exists no transfer function to the GW memory. These developments illustrate the evolving complexity of accurately modelling the LISA detector response in the frequency domain, particularly for complex signals such as those involving precession, eccentricity or GW memory. Furthermore, all these works assume an equilateral configuration of the LISA constellation.
@@
While the transfer function formalism shows promising results, it relies on several assumptions, like the equal arms configuration, which need to be validated against a more general formalism. The time domain approach can incorporate more realistic LISA configurations, providing a means to evaluate the robustness of the Fourier domain method. In this study, we take the first step in this direction by enabling the possibility of performing full LISA parameter estimation (PE) using time domain waveforms and responses. This is achieved by substantially reducing the computational cost of the likelihood evaluations for time domain signals.
@@
To achieve this, we will exploit the parallelization capabilities of phenomenological waveform models and employ a new implementation of the IMRPhenomT family Estellés et al. (2021, 2022a, 2022b) with GPU support, as well as the LISA response. We will conduct a series of signal injections, including aligned-spin and precessing systems with subdominant harmonics, to examine the accuracy with which LISA can recover source parameters. Additionally, we will investigate the influence of subdominant harmonics in the waveforms, the impact of omitting certain TDI channels, and the effect of using different TDI generations. These studies will provide key insights into the accuracy and efficiency of LISA parameter estimation methods.
@@ II.1 Waveform generation
The waveform generation, the calculation of the polarizations $h_{+,\times}(t)$, will be executed using a newly developed implementation of the IMRPhenomT family within the Python programming language with GPU support. Phenom models often utilize analytical expressions for describing the waveform as functions of time or frequency, where each data point is computed independently. This characteristic permits the parallelization of the waveform computation, a feature incorporated into this implementation through the use of Python libraries like numpy Harris et al. (2020) and numba Lam et al. (2015) for CPU operations, as well as cupy Okuta et al. (2017) for GPU acceleration. Importantly, this new implementation has been designed with agnostic code, allowing it to operate effectively on both CPU and GPU architectures. BBHx Katz et al. (2020); Katz (2022) is a similar software package designed for GPU implementation, supporting earlier generations of Fourier-based Phenom models, such as PhenomHM London et al. (2018), in addition to the Fourier LISA response.
The new implementation is designed to replicate the functionality found in the LIGO Algorithms Library (LALSuite) LIGO Scientific Collaboration et al. (2018); Wette (2020), but has been specifically tailored to extend the modularity and usability of phenomenological waveform models beyond the LIGO-Virgo-Kagra (LVK) infrastructure and its conventional sources.
@@
By facilitating a Python-based framework, this new package alleviates the challenges faced by waveform developers who traditionally need to reimplement their models in C99, the language supported by LALSuite. The adoption of Python allows for easier integration of advanced modeling techniques, such as machine learning and neural networks, which benefit from the more extensive library support available in Python.
@@
Moreover, the Phenom modelling strategy also facilitates the implementation of advanced algorithms that leverage non-uniform time and frequency grids, significantly reducing the number of required points. Notable methods include Multibanding Vinciguerra et al. (2017); García-Quirós et al. (2020); Morisaki (2021), heterodyne likelihood Cornish (2021), and relative binning Zackay et al. (2018); Leslie et al. (2021); Krishna et al. (2023). Efforts are also underway to adapt these algorithms for use in the time domain, further enhancing their applicability.
With this flexible implementation, we anticipate a significant boost in the development and adoption of phenomenological waveform models, which are crucial for meeting the computational efficiency and accuracy requirements, not only for LISA, but also for future third-generation ground-based detectors.
@@
The current implementation includes the complete IMRPhenomT family: IMRPhenomT Estellés et al. (2021) (aligned-spin, 22 mode only), IMRPhenomTHM Estellés et al. (2022a) (aligned-spin with subdominant modes), IMRPhenomTP Estellés et al. (2022b) (spin-precession with only the 22 mode in the co-precessing frame), IMRPhenomTPHM Estellés et al. (2022b) (spin-precession with co-precessing subdominant modes). In the precessing sector, the GPU acceleration is only fully implemented for the next-to-next leading order (NNLO) Bohé et al. (2013) prescription of the Euler angles describing the precessing motion. The Multi-scale Analysis (MSA) Chatziioannou et al. (2017) angles lack a special Bessel function that needs to be implemented in cupy. Finally, the numerical angles obtained from the evolution of the SpinTaylor equations Buonanno et al. (2004) can only use the GPU through linear interpolation of the solution in a coarser grid. Our analysis will only employ the NNLO analytical prescription with GPU support. By using this single-spin approximation, we will miss double-spin information and not measure very well the tilt and azimuthal angle of the secondary spin.
@@
Another notable advancement over the LALSuite implementation involves the precessing sector. The Phenom approach approximates precessional effects using the twisting-up technique Schmidt et al. (2012, 2015), which establishes a co-precessing frame that dynamically follows the motion of the orbital angular momentum (or the direction of maximum emission). This is achieved through a set of Euler angles, with various prescriptions available, such as the NNLO, MSA or SpinTaylor mentioned above. In the co-precessing frame, the waveform resembles an aligned spin one and is generated with the aligned spin model. Then it is transformed to an inertial frame by a time-dependent Euler angle rotation which incorporates the waveform modulations proper of precessing. In IMRPhenomT, two distinct rotations are employed: the first transforms the waveform from the co-precessing frame to the intermediate inertial $J$-frame, while the second transforms from the $J$-frame to the final $L_{0}$-frame which is the reference frame used in the LALSuite conventions. These two rotations and the calculation of the polarizations are formulated as
@@
In the new implementation, we utilize the efficient calculation of polarizations in rotating frames as detailed in Appendix B of Boyle et al. (2014), which we summarize in the following. By leveraging the relationship between the Wigner-D matrices and the spin-2 weighted spherical harmonics Boyle (2016) one obtains that the polarizations can be written as
@@
Thus, the computational complexity is reduced to a single summation over the number of modes in the co-precessing frame. The number of operations per time step is reduced to 10, 32, and 77 for the IMRPhenomT, a $l=5$ and a $l=8$ co-precessing model respectively. The Wigner-D matrices now encapsulate the resultant combinations of the two rotations between frames ($R_{CP-J}$, $R_{J-L_{0}}$) along with the alignment with the line-of-sight ($R_{\iota,\phi}$ [^1] The corresponding Euler angles for this last rotation are $\alpha=\phi$, $\beta=\iota$, $\gamma=0$.).
Consecutive rotations can also be represented as a unique rotation characterized by a new set of Euler angles. However, calculating these angles is highly complex. Each time step would require computing the rotation matrices from the angles (which involves trigonometric functions), performing the matrix multiplication and extracting the angles through inverse trigonometric functions. Besides being so computationally intensive, this method is also very sensitive to numerical errors [^2] For drawbacks of using Euler angles see Boyle ..
A more efficient approach for handling rotations involves using quaternions instead of Euler angles. Quaternions are four-dimensional constructs that obey specific rules of addition and multiplication, often used to represent rotations. The primary advantage is that successive rotations can be combined by simply multiplying quaternions, eliminating the need for complex matrix multiplications or costly inverse trigonometric functions at each time step. In this approach, we convert the Euler angles of each rotation into quaternions, then apply the quaternion multiplication to combine these rotations. The resulting quaternion is then used to compute the Wigner-D matrices. Given a quaternion $q=(q_{0},q_{1},q_{2},q_{3})$, its components are derived from the Euler angles as
@@
×$\times\sum_{\rho=0}^{l+m^{\prime}}\binom{\ell+m^{\prime}}{\rho}\binom{\ell-m^{\prime}}{\ell-\rho-m}\left(-\frac{|Q_{b}|^{2}}{|Q_{a}|^{2}}\right)^{\rho}.$
Figure 1: Mean evaluation time of a waveform generation (in Fourier domain) for different models and implementations as a function of total mass. Top pannel shows the aligned spin models while the bottom compares the precessing ones. The average is taken over 1000 randomly distributed cases with a frequency range of 0.1mHz-0.1Hz, therefore with a time sampling of 5s.
@@
In Fig. 1, we show the average computational cost of a waveform generation in the Fourier domain for LISA, comparing the performance of the new Python implementation with both the LALSuite-based implementation of IMRPhenonmT and IMRPhenomX families in LALSuite. For each model, we simulate 1000 random cases with mass ratios between [1-10] and total masses between $[10^{6}-10^{7}M\odot]$ logarithmically distributed. Spin parameters are uniformly distributed in the $z$-component for the aligned spin models and uniformly distributed in polar coordinates for the precessing models. We set the frequency range to 0.1 mHz–0.1 Hz with a sampling time of 5s. The total mass range is divided into 20 bins, and the average evaluation cost is computed within each bin.

### v3

@@ Abstract
We conduct the first full Bayesian inference analysis for LISA parameter estimation incorporating the effects of subdominant harmonics and spin-precession through a full time domain response. The substantial computational demands of using time domain waveforms for LISA are significantly mitigated by a novel Python implementation of the IMRPhenomT family of waveform models and the LISA response with GPU acceleration. This time domain response alleviates the theoretical necessity of developing specific transfer functions to approximate the LISA response in the Fourier domain for each specific type of systems and allows for the use of unequal-arms configurations and realistic LISA orbits. Our analysis includes a series of zero-noise injections for a Massive Black Hole Binary with aligned and precessing spins. We investigate the impact of including subdominant harmonics, compare equal and unequal-arm configurations, and analyze different Time-Delay-Interferometry (TDI) configurations. We utilize full and uniform priors, with a lower frequency cutoff of 0.1 mHz, and a signal duration of approximately two months, sampled every 5 seconds. The sampler is initialized based on Fisher’s estimates. Our results demonstrate LISA’s capability to measure the two spin magnitudes and the primary spin tilt angle, alongside sky localization, with percent-level precision, while component masses are determined with sub-percent accuracy.
@@ I Introduction
The Laser Interferometer Space Antenna (LISA) mission Colpi et al. (2024) was officially adopted by ESA and NASA in January 2024, with a scheduled launch for 2035. LISA will deploy three identical spacecraft into space to form the first space-based gravitational wave (GW) detector. Operating in space allows LISA to bypass many of the noise limitations that terrestrial detectors face Abbott et al. (2020), enabling the observation of a new class of GW sources. LISA will observe in the millihertz frequency range, detecting signals from a broad spectrum of sources, from stellar-mass compact objects to supermassive black holes, providing us with unprecedented insights into fundamental physics. LISA promises to be a transformative mission, set to revolutionize our understanding of the universe and GW astronomy.
@@
Moreover, LISA’s ability to observe a wide range of masses requires waveform models that encompass an expanded parameter space. At the same time, the high signal-to-noise ratios expected for LISA events will demand highly accurate models capable of capturing subtle physical effects that are less noticeable with current detectors. Key among these effects are precession and eccentricity, which provide crucial information about the astrophysical formation and evolutionary pathways of compact binaries Breivik et al. (2016); Steinle and Kesden (2021); Zevin et al. (2021); Fumagalli and Gerosa (2023); Fumagalli et al. (2024). Developing efficient parameter estimation techniques that incorporate both precession and eccentricity will be of paramount importance for fully exploiting LISA’s scientific potential.
@@
Parameter estimation for gravitational waves is typically performed in the Fourier domain due to the easier characterization of the noise and the efficiency and sensitivity of the matched filtering technique Allen et al. (2003, 2012), where the complicated convolutions and correlations between templates and data become a simple multiplication in the Fourier domain. Consequently, waveform models formulated directly in the Fourier domain Pratten et al. (2020); García-Quirós et al. (2020); Pratten et al. (2021); Cotesta et al. (2020) have gained significant popularity in GW data analysis Abbott et al. (2023); Colleoni et al. (2021); Mateu-Lucena et al. (2022). By avoiding the need for a discrete Fourier transform of each template, these models save valuable computational resources. In contrast, time domain waveforms require fine sampling on a uniformly spaced time grid to facilitate the application of the Fast Fourier Transform (FFT) algorithm. This constraint makes time domain waveforms particularly inefficient when dealing with long-duration signals, as the sampling grid must encompass the entire signal duration, even if one is interested only in a reduced frequency range. This also prevents the use of custom nonuniform frequency grids, which could be used to optimize performance Vinciguerra et al. (2017); García-Quirós et al. (2020); Canizares et al. (2015).
@@
However, from the modelling perspective, the time domain is often preferred because it naturally represents the temporal evolution of physical systems. Numerical Relativity (NR), Effective One Body formalism (EOB) Buonanno and Damour (1999); Damour and Nagar (2009) or post-Newtonian (PN) theory Will (2011); Blanchet (2014) solve the dynamic equations of motion in the time domain, yielding time domain waveforms that accurately describe the intricate behaviour of binary systems. In this domain, it is easier to incorporate physical processes and time-varying interactions, making it easier to asses how physical details like spin precession or eccentricity affect the waveform’s shape. Moreover, certain phenomena like displacement memory, a permanent change in spacetime following the passage of gravitational waves, are more naturally described and understood in the time domain Rosselló-Sastre et al. (2024). As LISA aims to explore new physics through these effects, time domain models will be crucial for maximizing the scientific return of the mission, enabling precise waveform modelling that is harder to replicate when relying solely on frequency domain representations.
@@
Apart from the challenges in waveform modelling, LISA introduces the additional complexity of computing a dynamic detector response. For current ground-based detectors, signals are typically short-lived, allowing the detector to be approximated as static. In these cases, the detector response is simply a linear combination of the two gravitational wave polarizations, with the response factors depending only on the sky position and polarization angle, but not on time. However, LISA’s sources can remain in-band for extended durations, from days to months or even years Colpi et al. (2024), causing the signal to be modulated by the orbital motion of the LISA constellation. Moreover, the laser frequency noise is several orders of magnitude larger than the GW signal LISA aims to detect. This noise is exacerbated when the LISA arms have unequal length. In a realistic scenario, the LISA spacecraft will not maintain a perfect equilateral triangular formation, leading to time-varying arm lengths and complicating gravitational wave detection. Fortunately, the Time-Delay Interferometry (TDI) technique Tinto and Dhurandhar (2014) mitigates this issue by combining signals from different arms in a carefully timed manner, effectively cancelling out laser frequency noise by several orders of magnitude. TDI channels reconstruct gravitational wave signals while suppressing noise, thereby enhancing the sensitivity of the detector to gravitational waves.
@@
The LISA arm response and TDI are better understood and described in the time domain. In this domain, they can be applied to the GW $h_{+,\times}(t)$ polarizations exhibiting any physical effect: higher harmonics, precession, eccentricity, or other complex features. However, applying these operations in the Fourier domain is more challenging and requires the development of theoretical transfer functions Marsat and Baker (2018); Marsat et al. (2021) to approximate the LISA response. These transfer functions rely on the assumption that a unique time-frequency correspondence exists in the gravitational wave signal. The requirement for a monotonic frequency and amplitude evolution complicates the Fourier domain approach, particularly for complex signals. Marsat et al. Marsat et al. (2021) addressed this issue for subdominant harmonics signals in aligned spin systems by developing a specific transfer function for each harmonic. Later, Pratten et al. Pratten et al. (2023) extended it to precessing systems by introducing the multiple-scale analysis to provide first-order corrections to the transfer function for each harmonic in precessing systems. These works marked significant progress, as they were the first to perform full Bayesian inference using complete inspiral-merger-ringdown (IMR) Fourier-domain waveforms for LISA. In contrast, Garg et al. Garg et al. (2024) applied this formalism to eccentric systems, using a Fourier-domain inspiral-only approximant without higher harmonics. So far, there exists no transfer function for the GW memory. These developments illustrate the evolving complexity of accurately modelling the LISA detector response in the frequency domain, particularly for complex signals such as those involving precession, eccentricity or GW memory. Furthermore, all these works assume an equilateral configuration of the LISA constellation.
@@
While the transfer function formalism shows promising results, it relies on several assumptions, like the equal-arms configuration, which need to be validated against a more general formalism. The time domain approach can incorporate more realistic LISA configurations, providing a means to evaluate the robustness of the Fourier domain method. In this study, we take the first step in this direction by enabling the possibility of performing full LISA parameter estimation (PE) using time domain waveforms and responses. This is achieved by substantially reducing the computational cost of the likelihood evaluations for time domain signals.
@@
To achieve this, we will exploit the parallelization capabilities of phenomenological waveform models and employ a new implementation of the IMRPhenomT family Estellés et al. (2021, 2022a, 2022b) and the LISA response with GPU support. We will conduct a series of signal injections, including aligned-spin and precessing systems with subdominant harmonics, to examine the accuracy with which LISA can recover source parameters. Additionally, we will investigate the influence of subdominant harmonics in the waveforms, the impact of unequal-arms orbits, the consequences of omitting certain TDI channels and the use of different TDI generations. These studies will provide key insights into the accuracy and efficiency of LISA parameter estimation methods.
@@ II.1 Waveform generation
The waveform generation, i.e. the calculation of the polarizations $h_{+,\times}(t)$, will be executed using a newly developed implementation of the IMRPhenomT family within the Python programming language with GPU support. Phenom models often utilize analytical expressions to describe the waveform as functions of time or frequency, where each time/frequency data point is computed independently from the rest. This characteristic permits the parallelization over the time/frequency array, a feature incorporated into this implementation through the use of Python libraries like numpy Harris et al. (2020) and numba Lam et al. (2015) for CPU operations, as well as cupy Okuta et al. (2017) for GPU acceleration. Importantly, this new implementation has been designed with agnostic code, allowing it to operate effectively on both CPU and GPU architectures. BBHx Katz et al. (2020); Katz (2022) is a similar software package designed for GPU implementation, supporting earlier generations of Fourier-based Phenom models, such as PhenomHM London et al. (2018), in addition to the Fourier LISA response.
@@
The new implementation is designed to replicate the functionality found in the LIGO Algorithms Library (LALSuite) LIGO Scientific Collaboration et al. (2018); Wette (2020), but has been specifically tailored to extend the modularity and usability of phenomenological waveform models beyond the LIGO-Virgo-Kagra (LVK) infrastructure and its conventional sources. By facilitating a Python-based framework, this new package alleviates the challenges faced by waveform developers who traditionally need to reimplement their models in C99, the language supported by LALSuite. The adoption of Python will allow for easier integration of advanced modeling techniques, such as machine learning and neural networks, which benefit from the more extensive library support available in Python.
@@
Moreover, the Phenom modelling strategy also facilitates the implementation of advanced algorithms that leverage nonuniform time and frequency grids, significantly reducing the number of required points. Notable methods include Multibanding Vinciguerra et al. (2017); García-Quirós et al. (2020); Morisaki (2021), heterodyne likelihood Cornish (2021), and relative binning Zackay et al. (2018); Leslie et al. (2021); Krishna et al. (2023). Efforts are also underway to adapt these algorithms for use in the time domain, further enhancing their applicability. With this flexible implementation, we anticipate a significant boost in the development and adoption of phenomenological waveform models, which are crucial to meet the computational efficiency and accuracy requirements, not only for LISA, but also for future third-generation ground-based detectors like Einstein Telescope Maggiore et al. (2020) and Cosmic Explorer Reitze et al. (2019).
@@
The current implementation includes the complete IMRPhenomT family: IMRPhenomT Estellés et al. (2021) (aligned-spin, (2,2) mode only), IMRPhenomTHM Estellés et al. (2022a) (aligned-spin with subdominant modes), IMRPhenomTP Estellés et al. (2022b) (spin-precession with only the (2,2) mode in the co-precessing frame), IMRPhenomTPHM Estellés et al. (2022b) (spin-precession with co-precessing subdominant modes). In the precessing sector, the GPU acceleration is only fully implemented for the next-to-next leading order (NNLO) Bohé et al. (2013) prescription of the Euler angles describing the precessing motion. The Multi-scale Analysis (MSA) angles Chatziioannou et al. (2017) will be incorporated soon. The numerical angles obtained from the evolution of the SpinTaylor equations Buonanno et al. (2004) can only use the GPU by interpolating or transfering the full ODE solution to the GPU. Our analysis will only employ the NNLO analytical prescription with GPU support. By using this single-spin approximation, we will miss double-spin information and not measure very well the tilt and azimuthal angle of the secondary spin.
@@
Another notable advancement over the LALSuite implementation involves the precessing sector. The Phenom approach approximates precessional effects using the twisting-up technique Schmidt et al. (2012, 2015), which establishes a co-precessing frame that dynamically follows the motion of the orbital angular momentum (or the direction of maximum emission). This is achieved through a set of Euler angles, with various prescriptions available, such as the NNLO, MSA or SpinTaylor mentioned above. In the co-precessing frame, the waveform resembles an aligned spin one and is generated with the aligned spin model. Then it is transformed to an inertial frame by a time-dependent Euler angle rotation which incorporates the waveform modulations proper of precession. In IMRPhenomT, two distinct rotations are employed: the first transforms the waveform from the co-precessing (CP) frame to the intermediate inertial $J$-frame, while the second transforms from the $J$-frame to the final $L_{0}$-frame which is the reference frame used in the LALSuite conventions. These two rotations and the calculation of the polarizations are formulated as
@@
In the new implementation, we utilize the efficient calculation of polarizations in rotating frames detailed in Appendix B of Boyle et al. (2014), which we summarize in the following. By leveraging the relationship between the Wigner-D matrices and the spin-2 weighted spherical harmonics Boyle (2016) one obtains that the polarizations can be written as
@@
Thus, the computational complexity is reduced to a single summation over the number of modes in the co-precessing frame. The number of operations per time step is reduced to 10, 32, and 77 for the IMRPhenomT, a $l=5$ and a $l=8$ co-precessing model respectively. The Wigner-D matrices now encapsulate the resultant combinations of the two rotations between frames $R_{CP-J}$, $R_{J-L_{0}}$ along with the alignment with the line-of-sight $R_{\iota,\phi}$ [^1] The corresponding Euler angles for this last rotation are $\alpha=\phi$, $\beta=\iota$, $\gamma=0$..
Consecutive rotations, as the ones considered here, can also be represented as a unique rotation characterized by a new set of Euler angles. However, calculating these new angles is highly complex. Each time step would require computing the rotation matrices from the Euler angles of each rotation (which involves trigonometric functions), performing the matrix multiplication and extracting the angles through inverse trigonometric functions. Besides being computationally intensive, this method is also very sensitive to numerical errors [^2] For drawbacks of using Euler angles see Boyle ..
A more efficient approach for handling rotations involves using quaternions instead of Euler angles. Quaternions are four-dimensional constructs that obey specific rules of addition and multiplication, often used to represent rotations. The primary advantage is that successive rotations can be combined by simply multiplying quaternions, eliminating the need for complex matrix multiplications or costly inverse trigonometric functions at each time step. In this approach, we convert the Euler angles of each rotation into quaternions, then apply the quaternion multiplication to combine these rotations. The resulting quaternion is then used to compute the Wigner-D matrices in Eq. 4. Given a quaternion $q=(q_{0},q_{1},q_{2},q_{3})$, its components are derived from the Euler angles as
@@
$\times\sum_{\rho=0}^{l+m^{\prime}}\binom{\ell+m^{\prime}}{\rho}\binom{\ell-m^{\prime}}{\ell-\rho-m}\left(-\frac{|Q_{b}|^{2}}{|Q_{a}|^{2}}\right)^{\rho}.$
@@
Figure 1: Mismatch distributions for the LALSuite and python implementations against NR SXS simulations. The aligned-spin model (IMRPhenomTHM) is matched against 1506 cases and the precessing version (IMRPhenomTPHM_NNLO) against 4152 in the mass range $10^{6},10^{7}]M_{\odot}$.
The new Python code also refines some aspects of the LALSuite implementation and enhances numerical stability by replacing finite-difference derivatives with analytical expressions, along with other improvements. We assess the accuracy of the new implementation by benchmarking it against numerical relativity waveforms from the LVCNR-SXS catalog Schmidt et al. (2017); Boyle et al. (2019) and compare the results to those obtained with the LALSuite implementation to ensure reliability. We evaluate 502 aligned-spin and 1384 precessing simulations from the SXS catalog Boyle et al. (2019). For each configuration, we generate three different cases by uniformly sampling the total mass in the range $10^{6},10^{7}]\,M_{\odot}$, the inclination angle from $0,\pi]$ radians, the reference phase $\phi_{\mathrm{ref}}$ from $0,2\pi]$ radians and the luminosity distance from $[1,10]$ Gpc. The minimum and reference frequencies are taken from the NR simulation metadata. In some cases, the simulations do not extend to sufficiently low frequencies (i.e., $10^{-4}$ Hz); in such cases, we compute the match starting from the minimum frequency available. We use a sampling interval of $\Delta t=5$ s and a frequency resolution of $\Delta f=10^{-6}$ Hz, corresponding to a time duration of approximately 11.6 days. The match is computed directly on the plus polarization using the sensitivity curve analytical_psd_lisa_tdi_AE as implemented in the PyCBC package Nitz et al. (2024). The mismatches are numerically optimized over the reference phase $\phi$ for the aligned-spin cases and over the reference phase and solid rotations of spins vectors for the precessing cases.
The results are summarized in Fig. 1. Our aim here is not to carry out a detailed mismatch study or to evaluate whether these models satisfy the LISA accuracy requirements, but rather to compare the mismatch values between the two implementations and verify that no significant differences arise. Consequently, the absolute mismatch values presented here should not be interpreted as statements about the overall waveform accuracy for LISA.
The results indicate that the mismatch distributions are nearly identical compared to NR, suggesting that the small improvements introduced in the Python implementation are negligible relative to the inherent modeling systematics with NR. Similar conclusions (not shown) are obtained by comparing with SEOBNRv5PHM Ramos-Buades et al. (2023) waveforms in both the LISA and LVK parameter regimes. We conclude that the new implementation is robust and suitable for use in this work. However, we stress that we are not claiming that these waveforms meet the full accuracy requirements of LISA. In the following section, we perform injections and recoveries using the same waveform model. This allows us to assess how well parameters could be recovered under the assumption of a perfect waveform model, with the understanding that this idealization does not hold in practice. The primary objective here is to demonstrate that parameter estimation using the full time-domain LISA response is computationally feasible, even for precessing binary systems.
Figure 2: Mean evaluation time of a waveform generation (in Fourier domain) for different models and implementations as a function of total mass. Top panel shows the aligned spin models while the bottom compares the precessing ones. The average is taken over 1000 randomly distributed cases with a frequency range of 0.1mHz-0.1Hz, therefore with a time sampling of 5s.
@@
In Fig. 2, we show the average computational cost of a waveform generation in the Fourier domain for LISA, comparing the performance of the new Python implementation with both the LALSuite-based implementation of IMRPhenonmT and IMRPhenomX families in LALSuite. For each model, we simulate 1000 random cases with mass ratios between [1-10] and total masses between $10^{6}-10^{7}M\odot]$ logarithmically distributed. Spin parameters are uniformly distributed in the $z$-component for the aligned spin models and uniformly distributed in polar coordinates for the precessing models. We set the frequency range to 0.1 mHz–0.1 Hz with a sampling time of 5 s. The total mass range is divided into 20 bins, and the average evaluation cost is computed within each bin.